OpenStudy (anonymous):

Please Simplify (2ab)^-2 over 2(ab)^3

OpenStudy (anonymous):

answer 2ab

jhonyy9 (jhonyy9):

1/(2ab)2/2(ab)3 = 2(ab)3/4(ab)2 = (ab)/2

OpenStudy (anonymous):

answer me this .... is (2ab) the same as 2(ab)?

OpenStudy (anonymous):

1/(2ab)2/2(ab)3 ----definitely agree with this bit, but don't how you came with 2(ab)3/4(ab)2 = (ab)/2

OpenStudy (anonymous):

yeah its ab/2

OpenStudy (anonymous):

not sure how you get from 1/(2ab)2/2(ab)3 to 2(ab)3/4(ab)2

jhonyy9 (jhonyy9):

check in my reply before on your next question